OpenStudy (anonymous):

dx/dt = 2 find dy/dt x= (-pi/3) y=tanx

OpenStudy (anonymous):

so its y=tan (-pi/3) find dy/dt

OpenStudy (anonymous):

\[y=\tan(x)\]\[\frac{dy}{dt}=\frac{dy}{dx}\frac{dx}{dt}\] \[\frac{dy}{dx}=\sec^2(x)\]\[\frac{dx}{dt}=2\]

OpenStudy (anonymous):

and so \(\frac{dy}{dt}=2\sec^2(x)\)